OpenStudy (anonymous):

2/3 divided 3/10 plz help

OpenStudy (bgrg007):

20/9

OpenStudy (anonymous):

how did you do it

OpenStudy (anonymous):

\[\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a}{b}*\frac{d}{c}\]

OpenStudy (bgrg007):